OpenStudy (anonymous):

3x-24=2X+1 i thought it was 25 but its not and i dont know what step i am missing

terenzreignz (terenzreignz):

Isn't it? Could you show your work?

OpenStudy (anonymous):



terenzreignz (terenzreignz):

And it's absolutely correct. I don't know why it wouldn't be :P

OpenStudy (anonymous):

the question says find the numeral value.. does that have something to do about it?

OpenStudy (anonymous):

some time given sol r wrong ,,,,

terenzreignz (terenzreignz):

does it say what the correct answer should be? (I still agree that it should be 25 though)

OpenStudy (anonymous):

it doesnt... it just keeps saying try again

terenzreignz (terenzreignz):

LOL report it :D We could be more sure though... @mathslover ? Your opinion, sensei? :)

OpenStudy (luigi0210):

Is this another angle problem?

OpenStudy (anonymous):

here is the question

OpenStudy (luigi0210):

Plug the number back in to find the length? :l

OpenStudy (luigi0210):

So 3x-24 3(25)-24 75-24 =51 So the length is 51 cm

OpenStudy (anonymous):

it said the answer was 51 any one know how they got that?

OpenStudy (luigi0210):

-_-'

mathslover (mathslover):

Yeah luigi is right. You dont' have to write the value of x in fact you have to find the value of 2x + 1 or 3x -24 3x -24 = 3(25) - 24 = 51 or 2x + 1 = 2(25) + 1 = 51

mathslover (mathslover):

Good work @Luigi0210 :)

OpenStudy (luigi0210):

Thank you @mathslover